OpenStudy (anonymous):

If f(x)=7x and g(x)=3x+1, find (f o g)(x). 21x^2+7x 21x+7 21x+1 10x+1

OpenStudy (anonymous):

i think its 21x+7 but im unsure @perl

OpenStudy (perl):

(f o g) (x) is a special symbol, it means f ( g ( x))

OpenStudy (perl):

lets calculate it

OpenStudy (anonymous):

ok great

OpenStudy (perl):

first plug in x into g(x) f ( g ( x ) ) = f ( 3x + 1 )

OpenStudy (perl):

when x goes into the g machine (g function) , out comes 3x + 1

OpenStudy (perl):

now we put 3x + 1 into the f machine as the new input

OpenStudy (anonymous):

ok heres where i usually get lost

OpenStudy (perl):

ok so 3x+1 is the new input for f now what is the rule for f machine?

OpenStudy (perl):

f machine says f(x) = 7*x treat 'x' as just a placeholder for input so what f machine really says is f ( input ) = 7 * input

OpenStudy (perl):

so using '3x+1' as your input you have f( 3x+1) = 7 * (3x + 1 )

OpenStudy (anonymous):

Thanks for the help @perl

OpenStudy (perl):

i like to draw the two functions as machines